OpenStudy (anonymous):

simplification of the expression (2b/c)^3

OpenStudy (anonymous):

2^3 b^3/ c^3 = 8b^3/c^3

OpenStudy (anonymous):

I LOVE YOU HAVE MY BABIESS

OpenStudy (unklerhaukus):

\[\left(\frac{xy}{z}\right)^n=\frac{x^ny^n}{z^n}\]